# RM-RL: Role-Model Reinforcement Learning for Precise Robot Manipulation

### v1

@@ Abstract
Precise robot manipulation is critical for fine-grained applications such as chemical and biological experiments, where even small errors (e.g., reagent spillage) can invalidate an entire task. Existing approaches often rely on pre-collected expert demonstrations and train policies via imitation learning (IL) or offline reinforcement learning (RL). However, obtaining high-quality demonstrations for precision tasks is difficult and time-consuming, while offline RL commonly suffers from distribution shifts and low data efficiency. We introduce a Role-Model Reinforcement Learning (RM-RL) framework that unifies online and offline training in real-world environments. The key idea is a role-model strategy that automatically generates labels for online training data using approximately optimal actions, eliminating the need for human demonstrations. RM-RL reformulates policy learning as supervised training, reducing instability from distribution mismatch and improving efficiency. A hybrid training scheme further leverages online role-model data for offline reuse, enhancing data efficiency through repeated sampling. Extensive experiments show that RM-RL converges faster and more stably than existing RL methods, yielding significant gains in real-world manipulation: 53% improvement in translation accuracy and 20% in rotation accuracy. Finally, we demonstrate the successful execution of a challenging task, precisely placing a cell plate onto a shelf, highlighting the framework’s effectiveness where prior methods fail.
I INTRODUCTION
@@
Autonomous robotic manipulation has shown promising results in real-world tasks, such as folding clothes and performing household manipulations [1, 2, 3]. Beyond these general-purpose applications, growing attention has shifted toward high-precision manipulation, particularly in delicate biological and chemical experiments [4, 5]. In such domains, success hinges on the robot’s ability to execute actions with sub-millimeter to millimeter-level accuracy, as even minor deviations can compromise the entire experiment.
@@
Two primary paradigms have been widely adopted for robotic policy learning: Imitation Learning (IL) and Reinforcement Learning (RL). Imitation Learning (IL) [6, 7] provides an effective way to acquire policies from expert demonstrations. However, in high-precision tasks, even human operators struggle to deliver consistent millimeter-level accuracy through teleoperation, making the collection of demonstrations slow and costly. Learning from such limited data often leads to overfitting to the training distribution and poor generalization in real-world scenarios. In contrast, Reinforcement Learning (RL) enables policies to autonomously explore and optimize through trial and error, thereby eliminating the need for demonstrations. Yet, most RL methods are developed and validated in simulators [8, 9, 10], which fail to capture the subtle but critical errors in fine-grained real-world manipulation due to the persistent sim-to-real gap. This motivates our goal: to enable efficient training of robotic policies with high-precision manipulation skills directly in the physical world.
@@
However, applying online RL directly in the real world is far from straightforward. In existing online RL, policies are typically updated directly from real-world reward signals. This process is highly data-inefficient, as each interaction can only be used once, and data collection in the physical world is inherently slow. Attempts to improve efficiency with replay buffers [11, 12, 13] partially mitigate this issue but introduce a second challenge: distribution shift [14, 15, 16]. Because stored data is generated by outdated policies, the mismatch between past experiences and the current policy often destabilizes optimization and impedes convergence. These limitations lead us to the central question of this work: Can we design a real-world RL framework that achieves both high data efficiency and training efficiency for precise robotic manipulation?
@@
“When three people walk together, there must be a role model whom I can learn from; I will select the good qualities and follow them.”   — Confucius, “The Analects”
To address this question, we draw inspiration from Confucius’ dictum in The Analects, and propose the Role-Model Reinforcement Learning (RM-RL) framework. RM-RL integrates online exploration with offline supervised learning by periodically selecting a role-model action, which is the highest-reward action observed under similar initial states, and using it to guide the remaining peer actions. The peer actions are automatically labeled by reference to the role-model action, allowing the online samples to be reformulated into supervised training data. These labeled samples are then repeatedly reused in offline training, ensuring that valuable experiences contribute to policy improvement multiple times rather than only once. This selection-and-labeling process is performed periodically during online interaction, which enables the framework to continuously inject stable supervised signals into the training loop. As a result, RM-RL substantially enhances data efficiency and mitigates distribution mismatch [17, 18], ultimately achieving both stable optimization and robust policy learning for high-precision robotic manipulation in the real world.
@@
II Related Works
II-A Reinforcement Learning
@@
Reinforcement learning (RL) aims to train a policy to make decisions based on the reward generated from the environment, typically modeled as a Markov decision process (MDP) [19]. Over the past decade, RL has demonstrated its power in combination with deep neural networks in multi-agent systems [20] and robotic areas, such as autonomous driving and robot locomotion tasks. However, RL still faces challenges when using real-world data for training [21].
@@
Traditional RL methods can be broadly categorized as online RL, where the policy model is updated simultaneously with the environment’s reward during the training process. Due to the emergence of high-performance simulators, like Isaaclab [10], Mujoco [9], online RL has demonstrated strong power in robotic locomotion, with Q value-based approaches (e.g., Q-learning and DQN [12]), policy gradient methods [22], and actor-critic frameworks (e.g., A3C [23], SAC [24]). Despite their success, online approaches often suffer from high sample complexity and safety concerns, making their deployment in real-world systems such as robotics and autonomous driving particularly challenging.
@@
To address the limitations of costly and risky online interaction, offline RL (also referred to as batch RL) has gained increasing attention. In offline RL, policies are trained entirely from previously collected datasets without additional environment access, allowing the reuse of experiences from demonstrations, simulators, or operational logs. Recent advances, including CQL [25], and IQL [26], have made notable progress toward improving stability and mitigating the risk of extrapolation errors, showing promising results in safety-critical domains such as healthcare, recommendation, and robot manipulation. Specifically, Zhou et al. [27] leverage an extensive dataset to improve the generalization capabilities. Nevertheless, offline RL still faces significant challenges in practice. Widely used offline RL datasets [28, 29, 30] are almost collected in simulation environments, where there exist significant gaps between the real-world settings. Collecting large-scale and high-quality real-world robotic datasets [31] remains difficult and resource-intensive, while the issue of distributional shift persists, as learned policies may generate out-of-distribution actions unsupported by the dataset. These limitations continue to hinder the scalability and reliability of offline RL in real-world applications.
II-B Reinforcement Learning for Real-world Robotic Tasks
Reinforcement Learning (RL) has demonstrated remarkable capabilities in robotic domains such as locomotion, manipulation, and autonomous navigation. Quadruped and humanoid robots have successfully employed deep RL to achieve agile locomotion directly on hardware [32, 33], while mobile robots have leveraged RL for navigation in unstructured environments [34, 35]. In manipulation, RL has enabled large-scale robotic grasping [36, 37, 38], underscoring its potential for practical deployment.
@@ II-B Reinforcement Learning for Real-world Robotic Tasks
Despite these successes, deploying RL in the real world remains challenging [21] due to issues of sample inefficiency, safety, and robustness to domain shifts. Collecting large amounts of on-robot data is expensive and risky, and simulators often fail to capture fine-grained physical properties, resulting in significant sim-to-real gaps [39, 40]. To address these challenges, researchers have developed approaches such as sim-to-real transfer with domain randomization [41], model-based RL with latent world models for efficient data usage [42], and offline RL leveraging prior robot datasets for safe policy training [26]. Moreover, hybrid frameworks that combine demonstrations with online fine-tuning [43] or integrate classical controllers with RL [44] further reduce the interaction burden and improve training stability.
III Methodology
@@
III-A Problem Definition
As we mentioned in Sec. I, our work addresses the robotic manipulating task requiring millimeter-level accuracy. The task involves using the robotic arm’s bio-gripper to pick and place a cell plate into a designated position of comparable size. We formalize the task as a standard one-step Markov Decision Process (MDP) [19], $\mathcal{M}=\big(\mathcal{S},\mathcal{A},p,\mathcal{R},\rho_{0},\gamma\big)$. Specifically, the state $\mathcal{S}$ includes the environmental images $I$ and the corresponding estimated picking poses $P^{e}$ of the robotic arm. The action $\mathcal{A}$ is the set of all candidate pose adjustments $\Delta P$ to $P^{e}$. For clarity, we denote each pose adjustment $\Delta P$ as an action $a\in\mathcal{A}$. The reward function $\mathcal{R}$ provides the real-world reward as the deviation between the final and target poses. Since the selected task is to pick and place the cell plate on a planar surface, we only consider the pose components $x$, $y$, and $\psi$, denoted as $P=[x,y,\psi]$. Here, $x$ and $y$ represent translations along the $x$- and $y$-axes, and $\psi$ denotes the rotation about the $z$-axis. Thus, the action can be simplified to $a=[\Delta x,\Delta y,\Delta\psi]$.
III-B Overall Framework
@@
Once the plate is successfully grasped, the robotic arm follows a pre-planned trajectory to place it at the center of the target area. The resulting pose $P^{final}_{i,t}$ is recorded, and the reward is computed by the reward function $\mathcal{R}$, based on the deviation between $P^{final}_{t}$ and $P^{target}$. This reward is used to update the policy network with the policy gradient algorithm [22]. In the Role-Model Online Labeling, we leverage the proposed role-model strategy to online label the training data from scene $i$ as $\mathcal{D}_{i}$, incrementally building a well-labeled dataset $\mathcal{D}$ for further offline RL training. The offline training processes are executed multiple times in the Role-Model Online Labeling and Data Replay. The details of the core contributions, role-model online labeling and online-offline training recipe, are presented in the following section.
III-C Role-Model Reinforcement Learning
@@
III-C1 Role-Model Online Labeling
@@
III-C2 Hybrid Online-Offline RL Training
@@
Online Training: The online RL adopts the policy gradient [45] fashion to train the policy network, since other training methods, like Actor-Critic [46, 47], Proximal Policy Optimization [48], require extensive action samples to achieve convergence, which is time-consuming in the real world. Online RL aims to maximize the following objective:
@@
Network Design: The architecture of our policy network $\pi_{\theta}$ is shown in Fig. 3. There are two inputs of the policy network, the color image $I$ and the estimated pose $P^{e}$. The output of the policy network is the adjustment pose $\Delta P$, referred to as the action $a\in\mathcal{A}$ introduced earlier, adjusting $P^{e}$ for precise picking. ResNet [49] is leveraged to extract the image feature, and an MLP is used to encode the input pose. Then, the image and pose features are concatenated and fed into three separate output heads, which respectively predict the translation adjustments along the $x$- and $y$-axes, and the rotation adjustment around the $z$-axis. All the heads adopt the same MLP architecture to predict logits, which are transformed into probability distributions using the softmax function. The adjustments $\Delta x$, $\Delta y$, and $\Delta\psi$, sampled from the predicted distributions, are represented collectively as the action $a=[\Delta x,\Delta y,\Delta\psi]$. Finally, the final picking pose is calculated as $P^{pick}=P^{e}+a$, which is executed by the robotic arm to pick the cell plate precisely.
@@
In summary, the Role-Model Reinforcement Learning (RM-RL) framework integrates online and offline training through a role-model strategy that provides labeled data for offline training. Unlike traditional RL with replay buffers [12], which suffers from distribution mismatch [11, 15], RM-RL constrains the policy network $\pi_{\theta}$ to move forward to the approximately optimal action among similar initial states by supervised learning, significantly improving sampling efficiency, adaptability, and stability.
III-D Cell Plate Localization
To get the estimated pose $P^{e}$ of the cell plate from the real-world observation, a general pipeline consisting of object detection, segmentation, and pose estimation is leveraged. For precisely localizing the cell plate, we use a global camera instead of some physical labels, providing both the environmental color image $I_{RGB}$ and depth image $I_{depth}$. The Grounding Dino [50], a pretrained open-vocabulary object detection method, is then utilized to localize the bounding box of the plate $(x_{min},y_{min},x_{max},y_{max})$ via the prompt “black cell plate”. Furthermore, we adopt Segment Anything Model (SAM) [51], a foundation model of object segmentation, to predict the mask area $\mathcal{M}$ from the bounding box of the target cell plate. The central pixel coordinate is denoted as $(u,v)$, where $u=\frac{x_{min}+x_{max}}{2}$ and $v=\frac{y_{min}+y_{max}}{2}$. Then, the depth value is the average depth within the cell plate mask, which can be calculated as:

### v2

@@ Abstract
Precise robot manipulation is critical for fine-grained applications such as chemical and biological experiments, where even small errors (e.g., reagent spillage) can invalidate an entire task. Existing approaches often rely on pre-collected expert demonstrations and train policies via imitation learning (IL) or offline reinforcement learning (RL). However, obtaining high-quality demonstrations for precision tasks is difficult and time-consuming, while offline RL commonly suffers from distribution shifts and low data efficiency. We introduce a Role-Model Reinforcement Learning (RM-RL) framework that unifies online and offline training in real-world environments. The key idea is a role-model strategy that automatically generates labels for online training data using approximately optimal actions, eliminating the need for human demonstrations. RM-RL reformulates policy learning as supervised training, reducing instability from distribution mismatch and improving efficiency. A hybrid training scheme further leverages online role-model data for offline reuse, enhancing data efficiency through repeated sampling. Extensive experiments show that RM-RL converges faster and more stably than existing RL methods, yielding significant gains in real-world manipulation: 53% improvement in translation accuracy and 20% in rotation accuracy. Finally, we demonstrate the successful execution of a challenging task, precisely placing a cell plate onto a shelf, highlighting the framework’s effectiveness where prior methods fail. Project site: https://ntumars.github.io/project/RMRL
[…]
Jianfei Yang at
1 INTRODUCTION
@@
Autonomous robotic manipulation has shown promising results in real-world tasks, such as folding clothes and performing household manipulations Zitkovich et al. (2023); Black et al. (2024); Kim et al. (2024). Beyond these general-purpose applications, growing attention has shifted toward high-precision manipulation, particularly in delicate biological and chemical experiments Lan et al. (2025); Li et al. (2024). In such domains, success hinges on the robot’s ability to execute actions with sub-millimeter to millimeter-level accuracy, as even minor deviations can compromise the entire experiment.
@@
Two primary paradigms have been widely adopted for robotic policy learning: Imitation Learning (IL) and Reinforcement Learning (RL). Imitation Learning (IL) Hussein et al. (2017); Ravichandar et al. (2020) provides an effective way to acquire policies from expert demonstrations. However, in high-precision tasks, even human operators struggle to deliver consistent millimeter-level accuracy through teleoperation, making the collection of demonstrations slow and costly. Learning from such limited data often leads to overfitting to the training distribution and poor generalization in real-world scenarios. In contrast, Reinforcement Learning (RL) enables policies to autonomously explore and optimize through trial and error, thereby eliminating the need for demonstrations. Yet, most RL methods are developed and validated in simulators Geng et al. (2025); Todorov et al. (2012); Mittal et al. (2023), which fail to capture the subtle but critical errors in fine-grained real-world manipulation due to the persistent sim-to-real gap. This motivates our goal: to enable efficient training of robotic policies with high-precision manipulation skills directly in the physical world.
@@
However, applying online RL directly in the real world is far from straightforward. In existing online RL, policies are typically updated directly from real-world reward signals. This process is highly data-inefficient, as each interaction can only be used once, and data collection in the physical world is inherently slow. Attempts to improve efficiency with replay buffers Levine et al. (2020); Mnih et al. (2015); Lillicrap et al. (2015) partially mitigate this issue but introduce a second challenge: distribution shift Rowland et al. (2019); Fujimoto et al. (2019); Liu et al. (2020b). Because stored data is generated by outdated policies, the mismatch between past experiences and the current policy often destabilizes optimization and impedes convergence. These limitations lead us to the central question of this work: Can we design a real-world RL framework that achieves both high data efficiency and training efficiency for precise robotic manipulation?
@@
“When three people walk together, there must be a role model whom I can learn from; I will select the good qualities and follow them."   — Confucius, “The Analects”
To address this question, we draw inspiration from Confucius’ dictum in The Analects, and propose the Role-Model Reinforcement Learning (RM-RL) framework. RM-RL integrates online exploration with offline supervised learning by periodically selecting a role-model action, which is the highest-reward action observed under similar initial states, and using it to guide the remaining peer actions. The peer actions are automatically labeled by reference to the role-model action, allowing the online samples to be reformulated into supervised training data. These labeled samples are then repeatedly reused in offline training, ensuring that valuable experiences contribute to policy improvement multiple times rather than only once. This selection-and-labeling process is performed periodically during online interaction, which enables the framework to continuously inject stable supervised signals into the training loop. As a result, RM-RL substantially enhances data efficiency and mitigates distribution mismatch Kumar et al. (2022); Sarker (2021), ultimately achieving both stable optimization and robust policy learning for high-precision robotic manipulation in the real world.
@@
2 Related Works
2.1 Reinforcement Learning
@@
Reinforcement learning (RL) aims to train a policy to make decisions based on the reward generated from the environment, typically modeled as a Markov decision process (MDP) Puterman (2014). Over the past decade, RL has demonstrated its power in combination with deep neural networks in multi-agent systems Busoniu et al. (2008) and robotic areas, such as autonomous driving and robot locomotion tasks. However, RL still faces challenges when using real-world data for training Dulac-Arnold et al. (2019).
@@
Traditional RL methods can be broadly categorized as online RL, where the policy model is updated simultaneously with the environment’s reward during the training process. Due to the emergence of high-performance simulators, like Isaaclab Mittal et al. (2023), Mujoco Todorov et al. (2012), online RL has demonstrated strong power in robotic locomotion, with Q value-based approaches (e.g., Q-learning and DQN Mnih et al. (2015)), policy gradient methods Williams (1992), and actor-critic frameworks (e.g., A3C Mnih et al. (2016), SAC Degris et al. (2012)). Despite their success, online approaches often suffer from high sample complexity and safety concerns, making their deployment in real-world systems such as robotics and autonomous driving particularly challenging.
@@
To address the limitations of costly and risky online interaction, offline RL (also referred to as batch RL) has gained increasing attention. In offline RL, policies are trained entirely from previously collected datasets without additional environment access, allowing the reuse of experiences from demonstrations, simulators, or operational logs. Recent advances, including CQL Kumar et al. (2020), and IQL Kostrikov et al. (2021), have made notable progress toward improving stability and mitigating the risk of extrapolation errors, showing promising results in safety-critical domains such as healthcare, recommendation, and robot manipulation. Specifically, Zhou et al. Zhou et al. (2022) leverage an extensive dataset to improve the generalization capabilities. Nevertheless, offline RL still faces significant challenges in practice. Widely used offline RL datasets James et al. (2020); Fu et al. (2020); Rafailov et al. (2024) are almost collected in simulation environments, where there exist significant gaps between the real-world settings. Collecting large-scale and high-quality real-world robotic datasets Walke et al. (2023) remains difficult and resource-intensive, while the issue of distributional shift persists, as learned policies may generate out-of-distribution actions unsupported by the dataset. These limitations continue to hinder the scalability and reliability of offline RL in real-world applications.
2.2 Reinforcement Learning for Real-world Robotic Tasks
@@
Reinforcement Learning (RL) has demonstrated remarkable capabilities in robotic domains such as locomotion, manipulation, and autonomous navigation. Quadruped and humanoid robots have successfully employed deep RL to achieve agile locomotion directly on hardware Hwangbo et al. (2019); Radosavovic et al. (2024), while mobile robots have leveraged RL for navigation in unstructured environments Wijayathunga et al. (2023); Liu et al. (2020a). In manipulation, RL has enabled large-scale robotic grasping Xu et al. (2023); Wan et al. (2023); Geng et al. (2023), underscoring its potential for practical deployment. Despite these successes, deploying RL in the real world remains challenging Dulac-Arnold et al. (2019) due to issues of sample inefficiency, safety, and robustness to domain shifts. Collecting large amounts of on-robot data is expensive and risky, and simulators often fail to capture fine-grained physical properties, resulting in significant sim-to-real gaps Zhao et al. (2020); Muratore et al. (2022). To address these challenges, researchers have developed approaches such as sim-to-real transfer with domain randomization Tobin et al. (2017), model-based RL with latent world models for efficient data usage Hafner et al. (2019), and offline RL leveraging prior robot datasets for safe policy training Kostrikov et al. (2021). Moreover, hybrid frameworks that combine demonstrations with online fine-tuning Nair et al. (2020) or integrate classical controllers with RL Johannink et al. (2019) further reduce the interaction burden and improve training stability.
3 Methodology
@@
3.1 Problem Definition
As we mentioned in Sec. 1, our work addresses the robotic manipulating task requiring millimeter-level accuracy. The task involves using the robotic arm’s bio-gripper to pick and place a cell plate into a designated position of comparable size. We formalize the task as a standard one-step Markov Decision Process (MDP) Puterman (2014), $\mathcal{M}=\big(\mathcal{S},\mathcal{A},p,\mathcal{R},\rho_{0},\gamma\big)$. Specifically, the state $\mathcal{S}$ includes the environmental images $I$ and the corresponding estimated picking poses $P^{e}$ of the robotic arm. The action $\mathcal{A}$ is the set of all candidate pose adjustments $\Delta P$ to $P^{e}$. For clarity, we denote each pose adjustment $\Delta P$ as an action $a\in\mathcal{A}$. The reward function $\mathcal{R}$ provides the real-world reward as the deviation between the final and target poses. Since the selected task is to pick and place the cell plate on a planar surface, we only consider the pose components $x$, $y$, and $\psi$, denoted as $P=[x,y,\psi]$. Here, $x$ and $y$ represent translations along the $x$- and $y$-axes, and $\psi$ denotes the rotation about the $z$-axis. Thus, the action can be simplified to $a=[\Delta x,\Delta y,\Delta\psi]$.
3.2 Overall Framework
@@
Once the plate is successfully grasped, the robotic arm follows a pre-planned trajectory to place it at the center of the target area. The resulting pose $P^{final}_{i,t}$ is recorded, and the reward is computed by the reward function $\mathcal{R}$, based on the deviation between $P^{final}_{t}$ and $P^{target}$. This reward is used to update the policy network with the policy gradient algorithm Williams (1992). In the Role-Model Online Labeling, we leverage the proposed role-model strategy to online label the training data from scene $i$ as $\mathcal{D}_{i}$, incrementally building a well-labeled dataset $\mathcal{D}$ for further offline RL training. The offline training processes are executed multiple times in the Role-Model Online Labeling and Data Replay. The details of the core contributions, role-model online labeling and online-offline training recipe, are presented in the following section.
3.3 Role-Model Reinforcement Learning
@@
3.3.1 Role-Model Online Labeling
@@
3.3.2 Hybrid Online-Offline RL Training
@@
Online Training: The online RL adopts the policy gradient Sutton et al. (1999) fashion to train the policy network, since other training methods, like Actor-Critic Grondman et al. (2012); Haarnoja et al. (2018), Proximal Policy Optimization Schulman et al. (2017), require extensive action samples to achieve convergence, which is time-consuming in the real world. Online RL aims to maximize the following objective:
@@
Network Design: The architecture of our policy network $\pi_{\theta}$ is shown in Fig. 3. There are two inputs of the policy network, the color image $I$ and the estimated pose $P^{e}$. The output of the policy network is the adjustment pose $\Delta P$, referred to as the action $a\in\mathcal{A}$ introduced earlier, adjusting $P^{e}$ for precise picking. ResNet He et al. (2016) is leveraged to extract the image feature, and an MLP is used to encode the input pose. Then, the image and pose features are concatenated and fed into three separate output heads, which respectively predict the translation adjustments along the $x$- and $y$-axes, and the rotation adjustment around the $z$-axis. All the heads adopt the same MLP architecture to predict logits, which are transformed into probability distributions using the softmax function. The adjustments $\Delta x$, $\Delta y$, and $\Delta\psi$, sampled from the predicted distributions, are represented collectively as the action $a=[\Delta x,\Delta y,\Delta\psi]$. Finally, the final picking pose is calculated as $P^{pick}=P^{e}+a$, which is executed by the robotic arm to pick the cell plate precisely.
@@
In summary, the Role-Model Reinforcement Learning (RM-RL) framework integrates online and offline training through a role-model strategy that provides labeled data for offline training. Unlike traditional RL with replay buffers Mnih et al. (2015), which suffers from distribution mismatch Levine et al. (2020); Fujimoto et al. (2019), RM-RL constrains the policy network $\pi_{\theta}$ to move forward to the approximately optimal action among similar initial states by supervised learning, significantly improving sampling efficiency, adaptability, and stability.
3.4 Cell Plate Localization
To get the estimated pose $P^{e}$ of the cell plate from the real-world observation, a general pipeline consisting of object detection, segmentation, and pose estimation is leveraged. For precisely localizing the cell plate, we use a global camera instead of some physical labels, providing both the environmental color image $I_{RGB}$ and depth image $I_{depth}$. The Grounding Dino Liu et al. (2024), a pretrained open-vocabulary object detection method, is then utilized to localize the bounding box of the plate $(x_{min},y_{min},x_{max},y_{max})$ via the prompt “black cell plate”. Furthermore, we adopt Segment Anything Model (SAM) Kirillov et al. (2023), a foundation model of object segmentation, to predict the mask area $\mathcal{M}$ from the bounding box of the target cell plate. The central pixel coordinate is denoted as $(u,v)$, where $u=\frac{x_{min}+x_{max}}{2}$ and $v=\frac{y_{min}+y_{max}}{2}$. Then, the depth value is the average depth within the cell plate mask, which can be calculated as: